How many years will it take to grow 675 to 1350 if the discount rate is i = 12% quarterly?
Answer #1

We use the formula:
A=P(1+r/4)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

1350=675*(1+0.12/4)^4n

(1350/675)=(1.03)^4n

Taking log on both sides;

log 2=4n*log 1.03

n=1/4[log 2/log 1.03]

which is equal to

=5.86 years(Approx).